OpenStudy (anonymous):

i have 3 eq A + B = 2 A + B + C = 4 AND A + C = 3 how do u find A B AND C?

OpenStudy (anonymous):

a=1 b=1 c=2

OpenStudy (anonymous):

A = 1 B = 1 C = 2 A and B have to both equal 1 or one equal 0 and the other equal 2. So that means 1+1+C=4, So C has to equal 2. You just have to make the equation make sense.

OpenStudy (anonymous):

2/2=1 so a and b are 1. then if you hav a b and c u know already a and b are 1 so a+b=2 then 4-2-=2 so now u know c is 2 so it will be easy to work out the next 1

OpenStudy (anonymous):

cheers